a chemist needs to mix 20L of a 40% acid solution with some 70% acid solution to obtain a mixture that is 50% acid. How many liters of the 70% acid solution should be used?
Answer by checkley77(12844)

.70x+.40*20=.50(20+x)
.70x+8=10+.50x
.70x-.50x=10-8
.20x=2
x=2/.2
x=10 L of 70% acid is used.
Proof:
.70*10+.40*.2=.50(20=10)
7+8=.5*30
15=15